# Imaginary gap-closed points and non-Hermitian dynamics in a class of dissipative systems

### v1

@@ IV.3 Model with varying loss rate $\gamma_{x}$
Figure 7: (a) Energy spectrums under PBCs. $t_{0}=0.3$ (red), $t_{0}=0.5$ (blue), $t_{0}=0.6$ (green). (b) The distribution of $P_{x}$ when $t_{0}=0.3$. (c), (d) The bulk distribution of $P_{x}$ in double logarithmic (c) and logarithmic (d) plots with non-uniform loss rates. For (b)-(d), system size $L=200$ and the initial position $x_{0}=150$. Throughout (a)-(d), $t_{p}=0.5$, $t_{1}=0.5$, $\gamma_{0}=0.01$, $\gamma_{c}=0.20$ are fixed.
@@
As shown in Fig. 7(a), this non-uniform system transitions from IGC to imaginary gaped when increasing the value of $t_{0}$. IGC points of the PBC spectrums also determine the bulk dynamics of quantum walk in this non-uniform system. When IGC points exist, $P_{x}$ follows a power-law decay in the bulk. Conversely, in the absence of points, $P_{x}$ follows an exponential law decay in the bulk. We can see this clearly in Fig. 7(c) and Fig. 7(d). Besides, a remarkable probability loss peak appeared at the edge when bulk $P_{x}$ shows a power-law decay, as shown in Fig. 7(b). This gives an explicit explanation of the edge burst phenomenon in Ref. [118].
@@ V Relation to the Liouvillian dynamics
Figure 8: (a) The corresponding bosonic quadratic open quantum system. (b), (c) The model only considers the nearest A-B coupling and chooses $\phi=\pi/2$, $t_{1}=0.5$, $t_{p}=0.5$, and $0.60.4<\gamma_{x}<$ 0.6 (randomly distributed). (b) Liouvillian spectrums. Blue: OBC, $t_{0}=0.3$; Red: PBC, $t_{0}=0.3$; Purple: OBC, $t_{0}=0.6$; Green: PBC, $t_{0}=0.6$. (c) The bulk distribution of $P_{x}$ in a double logarithmic plot.
When a quantum system is coupled to a Markovian bath, the corresponding dynamical evolution of the density matrix is governed by the master equation, which can be expressed in Lindblad formalism [21]
@@
where $H_{0}$ is the Hamiltonian of the system, $L_{\alpha}$ are the Lindblad dissipators describing quantum jumps randomly moving the state $\left|\psi\right\rangle$ to $L_{\alpha}\left|\psi\right\rangle$. Before the accident of a quantum jump, the short time evolution of the system is controlled by the effective Hamiltonian $\mathcal{H}_{eff}=H_{0}-i\sum_{\alpha}L_{\alpha}^{\dagger}L_{\alpha}$ as $d\rho/dt=i(\rho\mathcal{H}_{eff}^{\dagger}-\mathcal{H}_{eff}\rho)$.
In the open chain, we consider [Fig. 8(a)], the dissipators are given by $L_{x}=\sqrt{\gamma_{x}}c_{x}^{\mathtt{B}}$. Notably, the long time evolution of the system can be captured by the single correlation function $C_{xy}(t)=\mathtt{Tr}\left[\rho\left(t\right)c_{x}^{\dagger}c_{y}\right]$. For this dissipative system, any initial state will converge to the non-equilibrium steady state correlation $C(\infty)$, determined by $dC(\infty)/dt=0$. In this paper, we mainly consider the speed of converging to the steady state and focus on the deviation $\tilde{C}\left(t\right)=C\left(t\right)-C\left(\infty\right)$, whose evolution equation is given by
@@
Using a method similar to Sec. II, we can demonstrate that the damping matrix $X$ of the open chain under PBC also has universal Liouvillian gapless points. We can check our conclusion numerically with a specific model. Let the hopping $t_{m}=0,m=2,3,\cdots,n$, we can take the distribution of dissipators to be randomly distributed within 0.4 and 0.6 and get the Liouvillian spectrum [Fig. 8(b)]. We can see that if the condition of Eq. (12) can be satisfied, the Liouvillian spectrum can have Liouvillian gapless points, which correspond to the eigenmodes of the form $\left(1,0,e^{ik},0,e^{i2k},0,\cdots,e^{ikN},0\right)$. These modes can be viewed as dark modes, which lead to the algebraic convergence towards the steady state in the bulk [Fig. 8(c)]. Here, we already replaced the loss probability on B sites $P_{x}$ with the steady-state density on B sites $n^{B}_{x}$, which can be written as [120]

### v2

@@ IV.3 Model with varying loss rate $\gamma_{x}$
Figure 7: (a) Energy spectra under PBCs. $t_{0}=0.3$ (red), $t_{0}=0.5$ (blue), $t_{0}=0.6$ (green). (b) The distribution of $P_{x}$ when $t_{0}=0.3$. (c), (d) The bulk distribution of $P_{x}$ in double logarithmic (c) and logarithmic (d) plots with non-uniform loss rates. For (b)-(d), system size $L=200$ and the initial position $x_{0}=150$. Throughout (a)-(d), $t_{p}=0.5$, $t_{1}=0.5$, $\gamma_{0}=0.01$, $\gamma_{c}=0.20$ are fixed.
@@
As shown in Fig. 7(a), this non-uniform system transitions from IGC to imaginary gapped when increasing the value of $t_{0}$. IGC points of the PBC spectra also determine the bulk dynamics of quantum walk in this non-uniform system. When IGC points exist, $P_{x}$ follows a power-law decay in the bulk. Conversely, in the absence of points, $P_{x}$ follows an exponential law decay in the bulk. We can see this clearly in Fig. 7(c) and Fig. 7(d). Besides, a remarkable probability loss peak appeared at the edge when bulk $P_{x}$ shows a power-law decay, as shown in Fig. 7(b). This gives an explicit explanation of the edge burst phenomenon in Ref. [127].
@@ V Relation to the Liouvillian dynamics
Figure 8: (a) The corresponding bosonic quadratic open quantum system. (b), (c) The model only considers the nearest A-B coupling and chooses $\phi=\pi/2$, $t_{1}=0.5$, $t_{p}=0.5$, and $0.60.4<\gamma_{x}<$ 0.6 (randomly distributed). (b) Liouvillian spectra. Blue: OBC, $t_{0}=0.3$; Red: PBC, $t_{0}=0.3$; Purple: OBC, $t_{0}=0.6$; Green: PBC, $t_{0}=0.6$. (c) The bulk distribution of $P_{x}$ in a double logarithmic plot.
Leveraging the established correspondence between non-Hermitian dynamics and steady states of open quantum systems (as explored in Ref. [128]), our conclusions can be generalized to open quantum systems as well. When a quantum system is coupled to a Markovian bath, the corresponding dynamical evolution of the density matrix is governed by the master equation, which can be expressed in Lindblad formalism [21]
@@
where $H_{0}$ is the Hamiltonian of the system, $L_{\alpha}$ are the Lindblad dissipators describing quantum jumps randomly moving the state $\left|\psi\right\rangle$ to $L_{\alpha}\left|\psi\right\rangle$. In this system, the quantum dynamics started from single particle states are controlled by the effective Hamiltonian $\mathcal{H}_{eff}=H_{0}-i\sum_{\alpha}L_{\alpha}^{\dagger}L_{\alpha}$ as $d\rho/dt=i(\rho\mathcal{H}_{eff}^{\dagger}-\mathcal{H}_{eff}\rho)$.
In the open chain, we consider [Fig. 8(a)], the dissipators are given by $L_{x}=\sqrt{\gamma_{x}}c_{x}^{\mathtt{B}}$. Notably, the long time evolution of the system can be captured by the single-particle correlation function $C_{xy}(t)=\mathtt{Tr}\left[\rho\left(t\right)c_{x}^{\dagger}c_{y}\right]$. For this dissipative system, any initial state will converge to the non-equilibrium steady state correlation $C(\infty)$, determined by $dC(\infty)/dt=0$. In this paper, we mainly consider the speed of converging to the steady state and focus on the deviation $\tilde{C}\left(t\right)=C\left(t\right)-C\left(\infty\right)$, whose evolution equation is given by
@@
Using a method similar to Sec. II, we can demonstrate that the damping matrix $X$ of the open chain under PBC also has universal Liouvillian gapless points. We can check our conclusion numerically with a specific model. Let the hopping $t_{m}=0,m=2,3,\cdots$, we can take the distribution of dissipators to be randomly distributed, namely $L_{x}=\sqrt{\gamma_{x}}c_{x}^{B}$, where $\gamma_{x}$ is a random number within $0.4$ and $0.6$, it varies for each $x$. And we get the Liouvillian spectrum [Fig. 8(b)]. We can see that if the condition of Eq. (12) can be satisfied, the Liouvillian spectrum can have Liouvillian gapless points, which correspond to the eigenmodes of the form $\left(1,0,e^{ik},0,e^{i2k},0,\cdots,e^{ikN},0\right)$. These modes can be viewed as dark modes, which lead to the algebraic convergence towards the steady state in the bulk [Fig. 8(c)]. Here, we already replaced the loss probability on B sites $P_{x}$ with the steady-state density on B sites $n^{B}_{x}$, which can be written as [128]